# Mamba State-Space Models Are Lyapunov-Stable Learners

### v2

@@ 5.3 ICL is an emergent ability of instruction tuned Mamba SSMs
Instruction tuning greatly smooths ICL performance across both parameter classes. While instruction tuned SSMs and Transformers of 160 million parameters or fewer continue to display slight or detrimental ICL abilities, all parameters of 270 million and greater show positive ICL abilities. The instruction tuned OpenELM 450M model displays particularly impressive ICL abilities, but it is difficult to determine whether it is strictly due to architecture and/or pretraining recipe, or partially due to $143\%$ more total pretraining tokens than Mamba/Mamba-2 and Pythia models. For instruction tuned models of greater than 450 million parameters, all SSMs and Transformers show positive ICL abilities, with Mamba-2 2.7B greatly outperforming all other models (both SSM and Transformer) in this class.
Thus, in terms of ICL as a function of SSM model size, while no clear trend presents itself for pretrained models, ICL appears to emerge for instruction tuned Mamba and Mamba-2 SSMs of size 370 million and greater.
6 Discussion and Future Directions
Using dynamical systems theory, we’ve shown that the recurrent dynamics of Mamba SSMs are robust to small input perturbations. We’ve extensively confirmed this result, showing that: a) Mamba inference differences due to mixed-precision align with Transformers, (b) Mamba fine-tuning is significantly more robust to changes due to mixed-precision and PEFT than Transformers, and (c) combining MPFT and PEFT can more than halve training time and nearly triple memory efficiency for Mamba models. Using both MPFT and PEFT, we’ve shown that instruction tuning Mamba and Mamba-2 SSMs greatly narrows the pretraining ICL gap on natural language tasks relative to comparable Transformer LLMs. In particular, this allows Mamba-2 SSMs to greatly outperform the ICL abilities of a large number of instruction tuned, attention-based LLMs. Furthermore, complimentary to recent studies, we’ve shown that ICL for natural language tasks can be characterized as an emergent ability of Mamba and Mamba-2 models of 370 millions parameters or greater.
There are several avenues for future work. In particular, adapting Mamba’s CUDA kernels to support more aggressive low-precision PEFT methods [12] would further decrease the hardware needed to train Mamba models, while providing additional speedups and testing the limits of the derived stability results. Future studies could further explore what known emergent abilities Mamba SSMs are (and aren’t) capable of, as well as determining new Mamba-specific emergent abilities that are potentially not possible for Transformers to achieve. Furthermore, our theoretical contributions open the door for follow up studies, both in terms of extending our stability results to more general error (and adversarial) robustness results, as well as deriving new Mamba SSM-specific LoRA schemes which leverage the implicit weight tying achieved by targeting MambaBlock modules for either PEFT accuracy or efficiency improvements.
7 Acknowledgments
We thank Leidos for funding this research through the Office of Technology. Approved for public release 24-LEIDOS-0515-27826.
Appendix A Proof of weight-tying using LoRA in the MambaBlock
Due to the low-level nature of Mamba’s prefix scan optimizations (discussed in Section 3), standard use of LoRA adapters is made difficult within Mamba’s SSM-layer. E.g., while $B_{t},C_{t}$ and $\Delta_{t}$ are conceptually PyTorch linear layers, their bundling in a contiguous memory block and careful manipulation makes appending a LoRA adapter on any of these individual matrices non-trivial (particularly, while respecting the highly specialized layout of each LoRA adapters targeted layer). However, we note that the overall design of the MambaBlock’s hardware optimizations may be leveraged to both efficiently learn the parameter-space for the majority of time-varying parameters (thus achieving PEFT) and regularize parameters during training (thus improving fine-tuning generalization).
Theorem 1.
Consider the weight matrix $\mathbf{W}$ of a MambaBlock from Equation 4. Targeting $\mathbf{W}$ for LoRA during fine-tuning ties adaptation weights across $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$.
Proof.
Let $r$ be the specified LoRA dimension. Targeting this matrix for LoRA results in the adapter
$\tilde{\mathbf{W}}=$ $\mathbf{W}+\mathbf{W}^{\prime}$
$=$ $\mathbf{W}+\mathbf{U}\mathbf{V},$
where $\mathbf{U}\in\mathbb{R}{}^{n\times r}$, $\mathbf{V}\in\mathbb{R}{}^{r\times 3d}$, and $\mathbf{W}$ is frozen during fine-tuning. Thus, for index $[i,j]$,
$\mathbf{W}^{\prime}[i,j]=$ $\sum_{k=0}^{r-1}\mathbf{U}[i,k]\mathbf{V}[k,j].$
Recall the form of $\mathbf{W}$:
$\mathbf{W}[t-1,:d]=\texttt{diag}(\bm{\Delta}_{t}),\mathbf{W}[t-1,d:2d]=\texttt{diag}(\mathbf{B}_{t}),\mathbf{W}[t-1,2d:3d]=\texttt{diag}(\mathbf{C}_{t}),$
where $\mathbf{W}[0,:d]=\texttt{diag}(\bm{\Delta}_{1}),\mathbf{W}[n-1,d:2d]=\texttt{diag}(\mathbf{B}_{T})$, and so on. For index $[t-1,j]$, we thus have
$\tilde{\mathbf{W}}[t-1,j]=$ $\mathbf{W}[t-1,j]+\mathbf{W}^{\prime}[t-1,j]$
$=$ $\mathbf{W}[t-1,j]+\sum_{k=0}^{r-1}\mathbf{U}[t-1,k]\mathbf{V}[k,j].$
Thus, the weights $\mathbf{U}[t-1,:]$ are tied for any parameter $\tilde{\mathbf{W}}[t-1,j],j\in\{1,\dots,3d\}$, which are used to adapt parameters $\bm{\Delta}_{1},\mathbf{B}_{t},$ and $\mathbf{C}_{t}$.
Appendix B Mamba stable dynamics proof
@@
For input $(\bm{x}_{0},\mathbf{u}_{1})$ to a MambaBlock, small changes $(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time. That is, $\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
@@
$\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\lambda_{\mathtt{max}}{})}).$
@@
Theorem 2.
Let $(\bm{x}_{t-1},\mathbf{u}_{t})$ be the latent state and input at an arbitrary time $t\in[1,T]$ within a MambaBlock. Then small changes $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially decreasing over discrete-time, i.e., $\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
@@
Appendix C Mamba stable outputs proof
Theorem 3.
@@ Appendix C Mamba stable outputs proof
Consider $\varepsilon$-close inputs $(\bm{x}_{t-1},\mathbf{u}_{t})$ and $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$, and their respective outputs $\mathbf{y}_{t}$ and $\mathbf{y}_{t}^{\prime}$. Assume $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time. That is, $\max|F_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-F_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
@@
$\max|G_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-G_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|$ $\in\mathcal{O}(\varepsilon\exp{(N\zeta)})$
Appendix D Experimental Details
All model checkpoints were evaluated on all benchmarks and few-shot settings using the LM evaluation harness from Eleuther AI [16], version 0.4.2. Pythia and Mamba Huggingface checkpoints were used for all inference and fine-tuning experiments, e.g., EleutherAI/pythia-160m and state-spaces/mamba-130m-hf for the smallest respective models. All fine-tuning experiments were run using package versions Transformers 4.40.0.dev0, Accelerate 0.28.0, TRL 0.8.1, PyTorch 2.2.1+cu121, and PEFT 0.10.0. All Mamba-2 models were run using mamba-ssm v2.2.2 using Huggingface checkpoints, e.g., state-spaces/mamba-130m for the smallest model.
For MPFT, Flash Attention 2.0 [9] via flash_attn 2.5.7 was used for Pythia models. For FP16 and BF16 inference results, Flash Attention 2.0 was used for both Pythia and OLMo models. For OLMo results, the 336B-token checkpoint was used by specifying revision=step80000-tokens336B.
All Alpaca and OpenHermes fine-tuning experiments used the following training recipe (adapted from [50]): AdamW_torch optimizer, cosine annealing schedule, no gradient accumulation, maximum norm of 1.0 for gradient clipping, and no warmup steps. Training epochs used for all Alpaca and OpenHermes experiments were three and one, respectively. For both Pythia and Mamba models, the learning rate and LoRA dimension $r$ were scaled to improve performance of smaller models (per-model values listed in Table 2).
For SLL LoRA, targeted Mamba layers were {x_proj, embeddings, in_proj, out_proj}; x_proj is the large MambaBlock memory buffer which, when targeted by LoRA, regularizes the majority of SSM parameters during fine-tuning through weight tying (Theorem 3). Pythia targeted SLL LoRA layers were {dense, embed_in, query_key_value, dense_h_to_4h,dense_4h_to_h}, chosen to balance performance across model sizes.

### v3

@@ 4.4 Stable instruction tuning improves Mamba ICL for natural language tasks
Instruction tuning greatly smooths ICL performance across both parameter classes. Thus, while instruction tuned SSMs and Transformers of 160 million parameters or fewer continue to display slight or detrimental ICL abilities, all models of 270 million parameters and greater show positive ICL abilities, with Mamba-2 2.7B closely matching (or exceeding) the ICL ability of the top performing Transformer LLM per-shot. Furthermore, in terms of an emergent ability (Wei et al., 2022), ICL emerges for instruction tuned Mamba and Mamba-2 SSMs of size 370 million and greater, while no clear trend presents itself for pretrained models.
5 Conclusions and Future Directions
Using dynamical systems theory, we’ve shown that the recurrent dynamics of Mamba SSMs are stable to small input perturbations. We’ve extensively confirmed this result, showing that Mamba training is significantly more robust to changes introduced during MPFT and PEFT than Transformer-based LLMs across widely-used fine-tuning datasets and benchmarks. Furthermore, by including many more few-shot evaluations in our divergence metric than previous work, our stability studies uncovered that Transformer LLMs are susceptible to large deviation spikes during MPFT alone, as well as when MPFT is combined with PEFT. As MPFT and PEFT are two of the most common LLM fine-tuning frameworks, and as lower precision formats (e.g., NF4 and FP8) become more widely used for fine-tuning, we advocate for more extensive stability analysis (e.g., including several few-shot evaluations when calculating divergence) to expose large deviation spikes.
In order to mitigate large deviation spikes, we’ve shown that, unlike Transformer LLMs, Mamba models are extremely stable and do not exhibit this phenomena. Furthermore, in addition to stable MPFT and PEFT, we’ve shown that a core LLM reasoning ability (ICL) of Mamba models significantly improves with instruction tuning; i.e., we’ve shown that while pretrained Mamba models’ ICL capabilities lag behind those of pretrained LLMs on natural language tasks, instruction tuned Mamba models’ rival (and surpass, in many Mamba-2 cases) the ICL capabilities of instruction tuned Transformers. Thus, Mamba models are extremely stable learners with the potential to display the ICL capabilities of attention-based models once instruction tuned.
There are several avenues for future work. In particular, adapting Mamba’s CUDA kernels to support more aggressive low-precision PEFT methods (Dettmers et al., 2024) would further decrease the hardware needed to train Mamba models, while providing additional speedups and testing the limits of the derived stability results. Furthermore, our theoretical contributions open the door for several follow up studies. E.g., additional results building off Theorems 3 and  4 may tackle generalization error bounds and privacy fine-tuning (by considering a language modeling framework and calculating upper bounds under distributions of $\varepsilon$), and Mamba-specific decoding schemes (where new algorithms may exploit the derived deviation bounds ensure next-token predicitions lie within a deviation tolerance from the true optimal decoding).
6 Relations to Previous Work
MPFT divergence for LLMs. Previous work has studied the performance difference for attention-based LLMs under low-precision inference (Dettmers et al., 2022; Dettmers & Zettlemoyer, 2023) and fine-tuning (Dettmers et al., 2024). However, these works only measured divergence using single-shot performance (i.e., zero-shot in  (Dettmers et al., 2022; Dettmers & Zettlemoyer, 2023), 5-shot for MMLU and zero-shot otherwise in  (Dettmers et al., 2024)). In contrast, we report the difference in model performance as the mean divergence between the full-precision and mixed-precision models, averaged over the accuracy of four different few-shot evaluations. Thus, in contrast to previous work, our inclusion of multiple shots: a) uncovers significant divergence in widely used LLMs even at widely used mixed-precisions (FP16 and BF16), and (b) more rigorously tests mixed-precision training’s effect on one of LLMs’ most important reasoning abilities (ICL). We are not aware of any previous works which have attempted to understand the performance impact of MPFT on Mamba models.
Lyapunov stability and RNNs/SSMs. Lyapunov exponents have previously been considered (Mikhaeil et al., 2022; Vogt et al., 2022) for classic RNN structures (e.g., vanilla RNNs, LSTMs, GRUs, PLRNNs, etc.), to determine when such models exhibit chaotic dynamics and the impact on the exploding/vanishing gradient phenomena [^1] We note that this continues a long line of research exploring RNNs sensitivity to initial conditions and their subsequent ability to produce chaotic output (Ribeiro et al., 2020; Laurent & von Brecht, 2017; Bertschinger & Natschläger, 2004; Bertschinger et al., 2004), although previous work did not leverage Lyapunov exponents..
For S4 (Gu et al., 2022) SSMs, Goel et al. (2022) used Hurwitz matrices to characterize the numerical stability of linear time-invariant (LTI) S4 models. Similarly,  Orvieto et al. (2023) used the LTI property to derive the eigenvalue decomposition of S4-like models, providing stability conditions by bounding the corresponding eigenvalues (further discussed in Appendix E). However, such analysis is not applicable to time-varying models, such as Mamba, nor does it characterize the effects of sensitive dependence on initial conditions (e.g., divergence of two $\varepsilon$ close inputs). To the best of our knowledge, no previous works have used Lyapunov exponents to explore the effects of mixed-precision on recurrent neural models or Mamba architectures.
In-context learning. Recent work has sought to understand the in-context learning (ICL) capabilities of Mamba LLMs when trained from scratch for specific tasks (Park et al., 2024; Lee et al., 2024). Another line of recent work has sought to understand how hybrid Mamba-Transformer models may be directly distilled from Transformer models (Wang et al., 2024b). However, to the best of our knowledge, no existing works have either theoretically explored the effects small input changes (e.g., due to mixed-precision) have on Mamba’s recurrent dynamics or empirically explored such effects downstream impact on fine-tuning.
As previously noted, recent works (Park et al., 2024; Lee et al., 2024) have studied Mamba’s ability to perform ICL by training Mamba models for specific tasks. Such tasks include logistic regression, decision trees, and learning other simple function classes, following the work of Garg et al. (2022). We emphasize that, in this set up, relatively small Mamba models–33 million and 90 million parameters for Lee et al. (2024) and Park et al. (2024), respectively–are trained from scratch for every evaluated task. Indeed, Park et al. (2024) notes that subsequent work is necessary to understand Mamba’s ICL capabilities for language modeling using standard natural language benchmarks, as well as for larger model sizes. Thus, our study of both the pretrained and instruction tuned ICL capabilities of Mamba/Mamba-2 LLMs for natural language tasks are complimentary to previous works.
Failure Cases and Limitations
While we explored the use of LoRA for Mamba models, many other PEFT adapters exist (Liu et al., 2022; Li & Liang, 2021; Houlsby et al., 2019; Lester et al., 2021). Furthermore, while mixed-precision using FP16 and BF16 were explored, lower-precision methods exist (Dettmers et al., 2024). Both are interesting directions for future work. Finally, our timing and memory usage experiments using Alpaca did not consider the largest two Mamba models (1.4B and 2.8B) due to their exceeding A10G memory capacity for FP32 full fine-tuning.
Broader Impact Statement
The Mamba models considered are all LLMs, and thus have the same potential positive and negative societal impacts as other LLMs (e.g., hallucinations). Furthermore, fine-tuning is known to possibly erode existing LLM guardrails, and thus the described methods may be adapted for this fine-tuning use case (as is the case for all PEFT and MPFT methods). However, MPFT/PEFT is shown to improve the quality of Mamba models for downstream applications, which may be adapted for all positive LLM applications in society (e.g., personal assistants, task automation, code completion, etc.). Finally, the described methods decrease the computational constraints required to train and inference Mamba SSMs, which has implications for green ML (e.g., decreased CO2 emissions, positive climate change impact, etc.).
Acknowledgments
We thank Leidos for funding this research through the Office of Technology. This manuscript has been approved for public release 24-LEIDOS-0515-27826. We thank (NeurIPS) Reviewer Uruo, (TMLR) Reviewer z2sE, and (TMLR) Reviewer SpRc for helpful comments and feedback, which spurred the authors to derive and add several additional theoretical results. This work is dedicated to the memory of Bella Halloran; thank you for supporting this one final paper, you are missed.
Appendix A Mamba stable dynamics proof
@@
For input $(\bm{x}_{0},\mathbf{u}_{1})$ to a MambaBlock, small changes $(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time. That is, $\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(|\varepsilon|\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
@@
$\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(|\varepsilon|\exp{(N\lambda_{\mathtt{max}}{})}).$
@@
Theorem.
Let $(\bm{x}_{t-1},\mathbf{u}_{t})$ be the latent state and input at an arbitrary time $t\in[1,T]$ within a MambaBlock. Then small changes $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially decreasing over discrete-time, i.e., $\max|F_{\theta}^{N}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{N}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|\in\mathcal{O}(|\varepsilon|\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
@@
Appendix B Mamba stable outputs proof
Theorem.
@@ Appendix B Mamba stable outputs proof
Consider $\varepsilon$-close inputs $(\bm{x}_{t-1},\mathbf{u}_{t})$ and $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$, and their respective outputs $\mathbf{y}_{t}$ and $\mathbf{y}_{t}^{\prime}$. Assume $(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)$ produce deviations which are exponentially non-increasing over discrete-time. That is, $\max|F_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-F_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|\in\mathcal{O}(|\varepsilon|\exp{(N\zeta)})$, for some scalar $\zeta\leq 0$.
@@
$\max|G_{\theta}^{N}(\bm{x}_{t-1},\mathbf{u}_{t})-G_{\theta}^{N}(\bm{x}_{t-1}+\varepsilon,\mathbf{u}_{t}+\varepsilon)|$ $\in\mathcal{O}(|\varepsilon|\exp{(N\zeta)})$
Appendix C Mamba stable AMP and/or PEFT dynamics proof
For an arbitrary MambaBlock, let $\bm{x}_{0},\dots,\bm{x}_{T}$ and $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ be the trajectory of latent and input states under full-precision fine-tuning. During a forward pass, automatic mixed-precision (AMP) saves time and memory by computing forward activations in half-precision (FP16 or BF16). During a backward pass, AMP computes gradients in half-precision and up-casts to full-precision prior to updating. MFPT thus introduces changes to the inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{T}$ (which are passed through a Swish) and latent states $\bm{x}_{0},\dots,\bm{x}_{T}$ through $\bar{\bm{\Delta}}_{t}$ (which is passed through a softplus), up/downcasting of parameter updates, and the changes to the aforementioned changes to the input states.
Potentially in concert with MPFT, PEFT via LoRA induces changes by learning low-rank matrices $\hat{\mathbf{A}}_{t},\hat{\mathbf{B}}_{t},\hat{\mathbf{C}}_{t}$
$\hat{\bm{x}}_{t}$ $=(\bar{\mathbf{A}}_{t}+\hat{\mathbf{A}}_{t})\bm{x}_{t-1}+(\bar{\mathbf{B}}_{t}+\hat{\mathbf{B}}_{t})\mathbf{u}_{t}$
$\hat{\mathbf{y}}_{t}$ $=(\mathbf{C}_{t}+\hat{\mathbf{C}}_{t})\bm{x}_{t}$
Let $\bm{x}_{0}^{\prime},\dots,\bm{x}_{T}^{\prime}$ and $\mathbf{u}_{1}^{\prime},\dots,\mathbf{u}_{T}^{\prime}$ be the trajectory of latent and input states under MPFT and/or PEFT. We thus have the following
Theorem.
Let $\epsilon_{t}^{u}=\max|\mathbf{u}_{t}-\mathbf{u}_{t}^{\prime}|$, i.e., the maximum, positive-scalar difference between the $t$th MPFT and full fine-tuning input state. Similarly, let $\epsilon_{t}^{x}=\max|\bm{x}_{t}-\bm{x}_{t}^{\prime}|$ and $\epsilon^{*}=\max\bigcup_{t=0}^{T}\{\epsilon_{t}^{u},\epsilon_{t}^{x}\}$ (where we define $\mathbf{u}_{0}=0$). We thus have
$\max|F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{T}(\bm{x}_{0}^{\prime},\mathbf{u}_{1}^{\prime})|\in\mathcal{O}(\epsilon^{*}\exp{(T\zeta)}),$
for some scalar $\zeta\leq 0$. Thus, differences introduced by AMP and/or PEFT for Mamba models do not exponentially compound over discrete-time.
Proof.
Consider the $\bm{\epsilon}$-vector definition of the Lyapunov exponent: given $\lambda_{\mathtt{max}}{}$, $\bm{\epsilon}\in\mathbb{R}^{d}$, and two initial values $(\bm{x}_{0},\mathbf{u}_{1})$ and $(\bm{x}_{0}+\bm{\epsilon},\mathbf{u}_{1}+\bm{\epsilon})$, the maximum deviation between these points grows as:
$\max|F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{T}(\bm{x}_{0}+\bm{\epsilon},\mathbf{u}_{1}+\bm{\epsilon})|\in\mathcal{O}(\left\lVert\bm{\epsilon}\right\rVert_{2}\exp{(T\lambda_{\mathtt{max}}{})}).$
where $\left\lVert\right\rVert_{2}$ denotes the L2-norm for vectors.
Let $f(\cdot)$ be a function which maps a vector to the set of its elements, e.g., for $\bm{z}\in\mathbb{R}^{d},f(\bm{x})=\{x_{1},\dots,x_{d}\}$. Thus, note that for $a^{*}=\max\{|a|:\forall a\in f(\bm{\epsilon})\}$, $\left\lVert\bm{\epsilon}\right\rVert_{2}\leq a^{*}\sqrt{d}$ and $\mathcal{O}(\left\lVert\bm{\epsilon}\right\rVert_{2}\exp{(T\lambda_{\mathtt{max}}{})})\in\mathcal{O}(a^{*}\exp{(T\lambda_{\mathtt{max}}{})})$.
Let $\mathcal{E}=\bigcup_{t=0}^{T}\{\mathbf{u}_{t}-\mathbf{u}_{t}^{\prime},\bm{x}_{t}-\bm{x}_{t}^{\prime}\}$ (where we define $\mathbf{u}_{0}=0$), and $\epsilon^{*}=\max\{|a|:\forall a\in\bm{\epsilon},\forall\bm{\epsilon}\in\mathcal{E}\}$. $\forall\;\epsilon\in\mathcal{E}$, we thus have
$\max|F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{T}(\bm{x}_{0}+\epsilon,\mathbf{u}_{1}+\epsilon)|\in\mathcal{O}(\left\lVert\bm{\epsilon}\right\rVert_{2}\exp{(T\lambda_{\mathtt{max}}{})})\subseteq\mathcal{O}(\epsilon^{*}\exp{(T\lambda_{\mathtt{max}}{})}),$
Furthermore, $\{\mathbf{u}_{t}-\mathbf{u}_{t}^{\prime},\bm{x}_{t}-\bm{x}_{t}^{\prime}\}\subseteq\mathcal{E}$, and thus
$\max|F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{T}(\bm{x}_{0}-\bm{x}_{0},\mathbf{u}_{1}-\mathbf{u}_{1}^{\prime})|\in\mathcal{O}(\epsilon^{*}\exp{(T\lambda_{\mathtt{max}}{})})$
Let $\tau(t)$ be a function that maps time values such that $\tau(t)\in[1,T-t]$ and $\tau(t)=1,\tau(t+1)=2,\dots,\tau(t+T)=T-t$. Then $\mathbf{B}_{\tau(t)},\mathbf{C}_{\tau(t)},\bm{\Delta}_{\tau(t)}$ define a new MambaBlock with inputs $\mathbf{u}_{\tau(t)},\dots,\mathbf{u}_{\tau(t+T)}$ and subsequent recurrent states $\bm{x}_{\tau(t)-1},\bm{x}_{\tau(t)},\dots,\bm{x}_{\tau(t+T)}$. $\forall t\in\{1,\dots,T\}$, $\{\mathbf{u}_{t}-\mathbf{u}_{t}^{\prime},\bm{x}_{t}-\bm{x}_{t}^{\prime}\}\subseteq\mathcal{E}$, so that
$\max|F_{\theta}^{T-t}(\bm{x}_{\tau(t)}-1,\mathbf{u}_{\tau(t)})-F_{\theta}^{T-t}(\bm{x}_{\tau(t)-1}^{\prime},\mathbf{u}_{\tau(t)})|\in\mathcal{O}(\epsilon^{*}\exp{((T-t)\lambda_{\mathtt{max}}{})}).$
Thus, we have
$\max|F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{T}(\bm{x}_{0}^{\prime},\mathbf{u}_{1}^{\prime})|\in\mathcal{O}(\epsilon^{*}\exp{(T\lambda_{\mathtt{max}}{})}),$
where $\lambda_{\mathtt{max}}{}\leq 0$ by Lemma 1.
Appendix D MambaBlock Maximum Deviation results
D.1 $\varepsilon$ varying initial conditions
Figure 5: Perturbing the initial states of random MambaBlocks produces output state deviations which exponentially decrease over discrete time. $\mu$ denotes the maximum absolute deviation at discrete time $t$ averaged over all trials, while $\sigma$ denotes the respective standard deviation at each discrete time.
Figure 6: Perturbing the initial states of random MambaBlocks produces output state deviations which exponentially decrease over discrete time. $\mu$ denotes the maximum absolute deviation at discrete time $t$ averaged over all trials, while $\sigma$ denotes the respective standard deviation at each discrete time.
Appendix E Previous SSM stability results
Previous SSM stability results (Goel et al., 2022; Orvieto et al., 2023) consider the following linear time-invariant (LTI) equations used in S4 models,
$\bm{x}_{t}$ $=\bar{\mathbf{A}}\bm{x}_{t-1}+\bar{\mathbf{B}}\mathbf{u}_{t}$ (5)
$\mathbf{y}_{t}$ $=\mathbf{C}\bm{x}_{t}.,$ (6)
which can be unrolled to produce
$\bm{x}_{k}$ $=\sum_{j=0}^{k}\bar{\mathbf{A}}^{j}\bar{\mathbf{B}}\mathbf{u}_{j}.$ (7)
By eigendecomposition,
$\bar{\mathbf{A}}^{j}$ $=(\mathbf{P}\mathbf{\Lambda}\mathbf{P}^{-1})^{j}=\mathbf{P}\mathbf{\Lambda}^{j}\mathbf{P}^{-1}.$
In  (Orvieto et al., 2023), the derived sufficient condition for stability is thus $|\alpha_{i}|\leq 1,$ for $\mathbf{\Lambda}=\texttt{diag}([\alpha_{0},\alpha_{1},\dots,\alpha_{d}])$. Relatedly, in  (Goel et al., 2022), stability is enforced by constraining $\mathbf{\Lambda}$ to be Hurwitz, in which case the real part of each diagonal entry $\alpha_{i}$ is negative, i.e., $\operatorname{Re}(\alpha_{i})<0,$ for $\mathbf{\Lambda}=\texttt{diag}([\alpha_{0},\alpha_{1},\dots,\alpha_{d}])$.
Both results critically rely on the LTI property of Equations 5 and  6 to derive Equation 7. However, Mamba’s state-space equations are linear time-varying (LTV), and thus Equation 7 does not hold. Thus, previous SSM stability results do not directly apply to Mamba models.
Appendix F Experimental Details